OpenStudy (anonymous):

Solving Equations with Radicals. √x-2 =4

OpenStudy (campbell_st):

well square both sides of the equation, then add 2 that should give you the value of x if the equation is \[\sqrt{x -2} = 4\] if the equation is \[\sqrt{x} - 2 = 4\] add 2 to both sides of the equation then square both sides to find x hope it helps

OpenStudy (anonymous):

It;s the first one sorry I don;t know how to make that sign over all of them. (√x-2)^2 = (4)^2 x-2+2=16+2 x=18 Would this be how I would do it? @campbell_st

OpenStudy (campbell_st):

thats correct